# Data-dependent and Oracle Bounds on Forgetting in Continual Learning

### v2

@@ 4 Oracle bounds for forward and backward transfer
While the general upper bounds on backward transfer in Section 3 are useful for designing theoretically motivated algorithms (see Section 5), there is merit in trying to better understand the behavior of these bounds for specific posterior distributions. To that end, we consider bounds on performance relative to that of an oracle who knows the data-distribution, as opposed to the data-dependent bounds established in Section 3. Specifically, we consider the continual Gibbs learner $\hat{Q}^{\lambda_{t}}_{s:t}(h)=\frac{Q_{s}(h)e^{-\lambda_{t}\hat{\mathcal{L}}(h,S_{t})}}{\mathbb{E}_{h\sim Q_{s}}\left[e^{-\lambda_{t}\hat{\mathcal{L}}(h,S_{t})}\right]}~{}.$ The Gibbs learner is of particular interest in the context of analyzing bounds with KL-divergence as it provides an explicit expression for the divergence for any prior. Note that as opposed to the more general Gibbs posterior of (6), here the distribution takes into account $S_{t}$ in addition to $S_{s}$.
@@
For any $Q_{s},S_{s},\lambda_{t}>0$,
$\begin{split}\mathbb{E}_{S_{t}\sim\mathcal{D}_{t}}&\mathcal{L}(\hat{Q}^{\lambda_{t}}_{s:t},\mathcal{D}_{s})\leq\\ &\inf_{Q_{s:t}}\left\{\mathcal{L}(Q_{s:t},\mathcal{D}_{t})+\frac{1}{\lambda_{t}}D_{\mathrm{KL}}(Q_{s:t}||Q_{s})\right\}\\ &+\frac{\lambda_{t}K^{2}}{8m_{t}}+\frac{1}{\lambda_{t}}\log\mathbb{E}_{h\sim Q_{s}}\left[e^{\lambda_{t}\Delta\mathcal{L}(h,s,t)}\right].\end{split}$ (7)
Proof of Theorem 4.1 is provided in Appendix A. The main idea of the proof is to choose the Gibbs learner as the posterior in (LABEL:eq:forget-base2). Equation (7) contains a domain disagreement term, $\mathrm{Dis}(Q_{s},\mathcal{D}_{s},\mathcal{D}_{t},\lambda_{t})\triangleq\frac{1}{\lambda_{t}}\log\mathbb{E}_{h\sim Q_{s}}\left[e^{\lambda_{t}\Delta\mathcal{L}(h,s,t)}\right].$
Assuming tasks are sufficiently similar, we arrive at the following Corollary (proof in Appendix A).
Corollary 4.2.
For any given choice of $S_{s}$ and for any $\lambda_{t}>0$, if $\mathrm{Dis}(Q_{s},\mathcal{D}_{s},\mathcal{D}_{t},\lambda_{t})\leq\epsilon_{s,t}$, then $\mathbb{E}_{S_{s},S_{t}\sim\mathcal{D}_{s},\mathcal{D}_{t}}F(\hat{Q}^{\lambda_{t}}_{s:t},\mathcal{D}_{s})\leq\frac{\lambda_{t}K^{2}}{8m_{t}}+2\epsilon_{s,t}.$
4.1 Oracle bounds with discrepancy terms
As we can see from Theorem 4.1, the behavior of $Q_{s}$ with regard to both tasks can affect the bound significantly. While Theorem 4.1 demonstrates the importance of $Q_{s}$, its exact effect is unclear. Next we discuss several corollaries that provide a clearer role for $Q_{s}$ that defines its desired behavior. The basic oracle inequality theorem we start from is the following (proof in appendix A).
Corollary 4.3.
For any given $S_{s}\sim\mathcal{D}_{s},Q_{s},\lambda_{t}>0$,
$\mathbb{E}_{S_{t}\sim\mathcal{D}_{t}}$ $\mathcal{L}(\hat{Q}^{\lambda_{t}}_{s:t},\mathcal{D}_{s})\leq\frac{\lambda_{t}K^{2}}{8m_{t}}$
$+\frac{\mathbb{E}_{h\sim Q_{s}}\left[e^{\lambda_{t}\Delta\mathcal{L}(h,s,t)}\mathcal{L}(h,\mathcal{D}_{s})\right]}{\mathbb{E}_{h\sim Q_{s}}\left[e^{\lambda_{t}\Delta\mathcal{L}(h,s,t)}\right]},$ (8)
A more explicit bound can be derived from (4.3), see Corollary A.7 in the Appendix.
So far we have focused on oracle bounds for the Gibbs posterior for any prior $Q_{s}$. A further simplification can be obtained by considering the Gibbs prior (6), $Q_{s}=\hat{Q}^{\lambda_{t}}_{s}(h).$
From our definition of the domain disagreement term,
$\begin{split}\mathbb{E}_{S\sim\mathcal{D}_{s}}&\mathrm{Dis}(\hat{Q}^{\lambda_{t}}_{s},\mathcal{D}_{s},\mathcal{D}_{t},\lambda_{t})\leq\frac{\lambda_{t}K^{2}}{8m_{s}}-\mathcal{L}(P,\mathcal{D}_{t})\\ &-\frac{1}{\lambda_{t}}\mathbb{E}_{S\sim\mathcal{D}_{s}}\log\mathbb{E}_{h\sim P}\left[e^{-\lambda_{t}\hat{\mathcal{L}}(h,S)}\right].\end{split}$
Plugging this into (7) and using Jensen’s inequality, we obtain the following theorem.
@@ 4.1 Oracle bounds with discrepancy terms
For any $\lambda_{t}>0$, if $Q_{s}$ obeys (6),
$\begin{split}&\mathbb{E}_{S_{s}\sim\mathcal{D}_{s}}\mathbb{E}_{S_{t}\sim\mathcal{D}_{t}}\mathcal{L}(\hat{Q}^{\lambda_{t}}_{s:t},\mathcal{D}_{s})\leq\mathbb{E}_{S_{s}\sim\mathcal{D}_{s}}\mathcal{L}(\hat{Q}^{\lambda_{t}}_{s},\mathcal{D}_{t})\\ &+\frac{\lambda_{t}K^{2}}{8m_{t}}+\frac{\lambda_{t}K^{2}}{8m_{s}}+\mathcal{L}(P,\mathcal{D}_{s})-\mathcal{L}(P,\mathcal{D}_{t}).\end{split}$
This provides us with a more practical prior $\hat{Q}^{\lambda_{t}}_{s}$ at the cost of additional approximation error $\lambda_{t}K^{2}/8m_{s}$. If $m_{s},m_{t}\rightarrow\infty$ [^2] In this case, $\hat{Q}^{\lambda}_{s}=Q^{\lambda}_{s},\hat{Q}^{\lambda_{t}}_{s:t}=Q^{\lambda_{t}}_{s:t}$, we get an interesting bound involving generalization and forgetting,
$\begin{split}\mathbb{E}_{S_{s}\sim\mathcal{D}_{s}}\mathbb{E}_{S_{t}\sim\mathcal{D}_{t}}\mathcal{L}(\hat{Q}^{\lambda_{t}}_{s:t},\mathcal{D}_{s})&-\mathbb{E}_{S_{s}\sim\mathcal{D}_{s}}\mathcal{L}(\hat{Q}^{\lambda_{t}}_{s},\mathcal{D}_{t})\leq\\ &\mathcal{L}(P,\mathcal{D}_{s})-\mathcal{L}(P,\mathcal{D}_{t}).\end{split}$
We note that the right-hand-side is data-free and describes the general loss landscapes of both tasks with respect to the prior $P$. If both tasks come from the same distribution, for example, this implies that forgetting is upper bounded by the forward transfer (for Gibbs measures). This suggests that any problem with high forgetting for the Gibbs measure would also have poor forward transfer and a problem with good forward transfer will also have low forgetting.
4.2 Gibbs posteriors for correlated tasks
So far, we have derived oracle bounds that offer useful insights on the backward transfer for the Gibbs posterior. We would also like to examine whether stronger assumptions can offer bounds with improved or vanishing forgetting.
@@ 4.2 Gibbs posteriors for correlated tasks
For any $\lambda_{t}>0$, if (6) holds, with $\mathcal{L}(P,\mathcal{D}_{s},\mathcal{D}_{t})\triangleq\mathcal{L}(P,\mathcal{D}_{s})+\mathcal{L}(P,\mathcal{D}_{t})$,
$\mathbb{E}_{S_{s},S_{t}\sim\mathcal{D}_{s},\mathcal{D}_{t}}\mathcal{L}(\hat{Q}^{\lambda_{t}}_{s:t},\mathcal{D}_{s})\leq\frac{\lambda_{t}K^{2}}{8m_{t}}+\frac{\lambda_{t}K^{2}}{8m_{s}}$ (9)
$+\mathcal{L}(P,\mathcal{D}_{s},\mathcal{D}_{t})+\frac{1}{\lambda_{t}}\log\mathbb{E}_{h\sim P}\left[e^{-\lambda_{t}\mathcal{L}(h,\mathcal{D}_{t})}\right],$
$\mathbb{E}_{S_{s},S_{t}\sim\mathcal{D}_{s},\mathcal{D}_{t}}\mathcal{L}(\hat{Q}^{\lambda_{t}}_{s:t},\mathcal{D}_{t})\leq\frac{\lambda_{t}K^{2}}{8m_{t}}+\frac{\lambda_{t}K^{2}}{8m_{s}}$
$+\mathcal{L}(P,\mathcal{D}_{s},\mathcal{D}_{t})+\frac{1}{\lambda_{t}}\log\mathbb{E}_{h\sim P}\left[e^{-\lambda_{t}\mathcal{L}(h,\mathcal{D}_{s})}\right].$
We can see that in both cases, the loss is bounded by functions of sample size originating from the deviations of sample mean from its true mean (e.g., Hoeffding’s inequality), and the loss of the data-free prior, which serves as a limit if no additional information on the tasks is available.
One such source of additional information is the loss covariance, measuring task similarity,
We note that this covariance term is bounded in $[-1,1]$.
Corollary 4.6.
For any $\lambda>0$, if (6) holds, and $\mathrm{cov}_{\lambda_{t}}(P,s,t)\geq 0$,
$\mathbb{E}_{S_{s},S_{t}\sim\mathcal{D}_{s},\mathcal{D}_{t}}\mathcal{L}(\hat{Q}^{\lambda_{t}}_{s:t},\mathcal{D}_{s})\leq\frac{\lambda_{t}K^{2}}{8m_{s}}+\mathcal{L}(P,\mathcal{D}_{s})$
$\mathbb{E}_{S_{s},S_{t}\sim\mathcal{D}_{s}}\mathcal{L}(\hat{Q}^{\lambda_{t}}_{s:t},\mathcal{D}_{t})\leq\frac{\lambda_{t}K^{2}}{8m_{t}}+\mathcal{L}(P,\mathcal{D}_{t}).$
We can further improve upon this bound given a tighter bound on the covariance.
Corollary 4.7.
For any $\lambda>0$, if (6) holds, and $\mathrm{cov}_{\lambda_{t}}(P,s,t)\geq e^{-c}$,
$\mathbb{E}_{S_{s},S_{t}\sim\mathcal{D}_{s},\mathcal{D}_{t}}\mathcal{L}(\hat{Q}^{\lambda_{t}}_{s:t},\mathcal{D}_{s})\leq\frac{\lambda_{t}K^{2}}{8m_{s}}+\frac{c}{\lambda_{t}}$
$\mathbb{E}_{S_{s},S_{t}\sim\mathcal{D}_{s},\mathcal{D}_{t}}\mathcal{L}(\hat{Q}^{\lambda_{t}}_{s:t},\mathcal{D}_{t})\leq\frac{\lambda_{t}K^{2}}{8m_{t}}+\frac{c}{\lambda_{t}}.$
Both Corollaries 4.6 and 4.7 are specific cases of more general Theorems that apply for any finite number of tasks. These bounds will be presented in the following subsection.
We observe that if the covariance is negative, Corollaries 4.6 and 4.7 do not hold, and we cannot improve on (9).This is unsurprising, as it means that the hypotheses that perform well on the source task perform poorly on the target and vice versa, which makes learning to generalize without forgetting impossible unless the initial prior is somehow already near-optimal. We note that the bound in Theorem 4.6 does not converge to zero as sample size goes to infinity, and the final term is in fact independent of the sample size. This is due to the fact that the covariance being non-negative is a necessary condition to guarantee that the backward transfer does not increase. As we can see in Corollary 4.7, a higher covariance leads to bounds that converge to zero upon appropriately choosing $\lambda_{t}$, as sample size increases.
Note that the phenomenon of self-forgetting seen in the NTK regime in (Karakida and Akaho, 2021) can also be seen here: if $m_{s}$ is small we can get forgetting even if the target task is identical (high covariance), and learning the same task again with new data may be worse than using all of the data as a single training set.
4.3 Extension to $T$ tasks
Suppose we are given a set of tasks $\{1,2,\ldots,T\}=[T]$, appearing sequentially. We would like to minimize the average backward transfer after all tasks, $BWT(Q_{1:T},[T])$. In general, this definition does not assume anything about the construction of the posterior, other than the fact that we have no access to samples from future (unseen) tasks while learning each specific posterior $Q_{1:t}$.
Starting from (3.3), we show in appendix A that if all of the priors are empirical Gibbs distributions, namely $\forall i\in\{2,\ldots,T\},~{}~{}\hat{Q}^{\lambda_{i}}_{1:i}(h)\propto\hat{Q}^{\lambda_{i-1}}_{1:i-1}(h)e^{-\lambda_{i}\hat{\mathcal{L}}(h,S_{i})}$, where $\hat{Q}^{\lambda_{1}}_{1:1}(h)\propto P(h)e^{-\lambda_{1}\hat{\mathcal{L}}(h,S_{1})}$, we have the following oracle bound.
Theorem 4.8.
$\mathrm{(Forgetting)}$ For any $\lambda_{T}>0$, assuming all $Q_{1:j}$ are empirical Gibbs posteriors, and that $\mathrm{cov}_{P}(i,[T])\geq 0$, for any sample of training sets $S_{j}\sim\mathcal{D}_{j}$, $\forall i\in[T-1]$
@@ 4.3 Extension to $T$ tasks
The main ideas of the proof are to make use of the structure of Gibbs distributions to arrive at an explicit term for the KL-divergence, then make use of the assumption on the covariance to decompose said term into the loss on task $i$ and all other losses. As the extension to the covariance for two tasks, we have a similar notion of covariance
As far as forward transfer, we can apply a similar analysis (for standard change-of-measure).
Theorem 4.9.
$\mathrm{(Transfer)}$ For any $\lambda_{T}>0$, if all $Q_{1:j}$ are empirical Gibbs posteriors, and $\mathrm{cov}_{P}(T,[T])\geq 0$, for any sample of training sets $S_{j<T}\sim\mathcal{D}_{j}$,
$\begin{split}\mathbb{E}_{S_{T}}\mathcal{L}(\hat{Q}^{\lambda_{T}}_{1:T},\mathcal{D}_{T})&\leq\frac{\lambda_{T}K^{2}}{8m_{T}}+\mathcal{L}(P,\mathcal{D}_{T}).\end{split}$ (11)
Theorem 4.8 is somewhat surprising, as it implies a sufficient condition for learning without forgetting that, for each individual task, does not become worse with the number of tasks and has no direct dependence on task order or on the length of time a task has not been seen. While the r.h.s. in (10) contains a constant term $\mathcal{L}(P,\mathcal{D}_{i})$, this term does not depend on the number of total tasks $T$ or on $i$. We also note that the forgetting $F(\hat{Q}^{\lambda_{T}}_{1:T},[T])$ may be negative. We show in Corollary A.9 that a stronger covariance condition even leads to vanishing forgetting, similarly to Corollary 4.7.
This lack of dependence on task order can be attributed to the nature of the empirical Gibbs distribution that applies a combination of exponential weights on the initial distribution $P$. The final weight of any given hypothesis $h$ therefore does not depend on the order of tasks but rather only on $P(h)$ and its empirical performance on each task.
The fact that this bound does not degrade with the number of tasks is a result of the assumption on the non-negative covariance: we assume that hypotheses that perform well on task $i$ tend to perform well on all other tasks, and thus the exponential weighting scheme does not significantly reduce their probability in the final distribution. While this is a somewhat strong assumption, it is weaker than assuming that a single optimal solution to all tasks exists.

### v3

@@ 4 Oracle bounds for backward transfer
While the general upper bounds on backward transfer in Section 3 are useful in practice, as well as for designing theoretically motivated algorithms (see Section 5), there is merit in trying to better understand the behavior of these bounds for specific posterior distributions. To that end, we consider bounds on performance relative to that of an oracle who knows the data-distribution, as opposed to the data-dependent bounds established in Section 3. Specifically, we consider the continual Gibbs learner $\hat{Q}^{\lambda_{t}}_{s:t}(h)=\frac{Q_{s}(h)e^{-\lambda_{t}\hat{\mathcal{L}}(h,S_{t})}}{\mathbb{E}_{h\sim Q_{s}}\left[e^{-\lambda_{t}\hat{\mathcal{L}}(h,S_{t})}\right]}~.$ The Gibbs learner is of particular interest in the context of analyzing bounds with KL-divergence as it provides an explicit expression for the divergence for any prior, as well as being the optimal posterior for minimizing the r.h.s. in Theorem 3.1. In addition, we note that if $\lambda\rightarrow\infty$, the Gibbs posterior converges to the empirical risk minimizer (ERM), meaning we can potentially apply our upper bounds to more traditional methods such as SGD.
Corollary 4.1.
For any $Q_{s},S_{s},\lambda_{t}>0$,
$\begin{split}\mathbb{E}_{S_{t}\sim\mathcal{D}_{t}}&\mathcal{L}(\hat{Q}^{\lambda_{t}}_{s:t},\mathcal{D}_{s})\leq\inf_{Q_{s:t}}\left\{\mathcal{L}(Q_{s:t},\mathcal{D}_{t})+\frac{1}{\lambda_{t}}D_{\mathrm{KL}}(Q_{s:t}||Q_{s})\right\}+\frac{\lambda_{t}K^{2}}{8m_{t}}+\frac{1}{\lambda_{t}}\log\mathbb{E}_{h\sim Q_{s}}\left[e^{\lambda_{t}\Delta\mathcal{L}(h,s,t)}\right].\end{split}$ (4)
Proof of Corollary 4.1 is provided in Appendix B. The main idea of the proof is to choose the Gibbs learner as the posterior in (LABEL:eq:forget-base2). As we can see from Corollary 4.1, the behavior of $Q_{s}$ with regard to both tasks can affect the bound significantly, though its exact effect is unclear.
In order to better understand the overall behavior of the bound’s elements, as well as examining whether stronger assumptions can offer more insightful bounds, we define a measure of the task similarity based on the loss covariance between the tasks:
We note that this covariance term is bounded in [1,1][-1,1].
Corollary 4.2.
For any $\lambda_{t}>0$, if (3) holds, and $\mathrm{cov}_{\lambda_{t}}(P,s,t)\geq 0$,
$\mathbb{E}_{S_{s},S_{t}\sim\mathcal{D}_{s},\mathcal{D}_{t}}\mathcal{L}(\hat{Q}^{\lambda_{t}}_{s:t},\mathcal{D}_{s})\leq\frac{\lambda_{t}K^{2}}{8m_{s}}+\mathcal{L}(P,\mathcal{D}_{s})\quad;\quad\mathbb{E}_{S_{s},S_{t}\sim\mathcal{D}_{s}}\mathcal{L}(\hat{Q}^{\lambda_{t}}_{s:t},\mathcal{D}_{t})\leq\frac{\lambda_{t}K^{2}}{8m_{t}}+\mathcal{L}(P,\mathcal{D}_{t}).$
This Corollary is a specific case of the more general Theorem 4.5 that applies for any finite number of tasks. Like the more general Theorem, the loss covariance term depends on the expressive power of the hypothesis class $\mathcal{H}$ as well as the similarity between the tasks.
4.1 Oracle bounds for multiple tasks
As in previous sections, we would like to extend our results to more than two tasks, as well as attempt to derive bounds that are easier to quantify. In this subsection, we will derive an asymptotic oracle upper bound on the backward transfer loss for the empirical Gibbs posterior.
Suppose we are given a set of tasks $T]\triangleq\{1,2,\ldots,T\}$, appearing sequentially. For simplicity we will assume that each task has identical sample of size $m$. We would like to minimize the average backward transfer after all tasks, $BWT(Q_{1:T},[T])$. In general, this definition does not assume anything about the construction of the posterior, other than the fact that we have no access to samples from future (unseen) tasks while learning each specific posterior $Q_{1:t}$.
In order to obtain quantifiable asymptotic bounds on the backward transfer, we make several simplifying assumptions on our hypothesis space and loss function:
Assumption 2.
The following five claims hold: (1) The hypothesis space $\mathcal{H}$ is a compact bounded subset of $\mathbb{R}^{d}$. (2) $\ell(h,z)$ is twice continuously differentiable w.r.t. $h$. (3) For all $i\in[T-1]$, the function $\sum_{t=1,t\neq i}^{T}\mathcal{L}(h,\mathcal{D}_{t})$ has a unique global minimum marked $h^{*}_{1:T/i}$. (4) For all $i\in[T-1]$, $P(h^{*}_{1:T/i})>0$. (5) For $m\rightarrow\infty$  [^2] This holds if data within each task is sampled i.i.d.
$\mathbb{E}_{S_{1},\ldots,S_{T}}\mathrm{cov}_{P}\left(e^{\lambda_{T}\sum_{j=1}^{T}\mathcal{L}(h,\mathcal{D}_{j})},e^{-\lambda_{T}\sum_{j=1}^{T}\hat{\mathcal{L}}(h,S_{j})}\right)\leq 0.$
Starting from (4), we show in appendix B that if all of the priors are empirical Gibbs distributions, namely $\forall i\in\{2,\ldots,T\},~~\hat{Q}^{\lambda_{i}}_{1:i}(h)\propto\hat{Q}^{\lambda_{i-1}}_{1:i-1}(h)e^{-\lambda_{i}\hat{\mathcal{L}}(h,S_{i})}$, where $\hat{Q}^{\lambda_{1}}_{1:1}(h)\propto P(h)e^{-\lambda_{1}\hat{\mathcal{L}}(h,S_{1})}$, we have the following asymptotic oracle bound.
@@ 4.1 Oracle bounds for multiple tasks
For any $\lambda_{T}>0$, assuming all $Q_{1:j}$ are empirical Gibbs posteriors, and that Assumption 2 holds, the following holds a.s. in expectation over $S_{j}\sim\mathcal{D}_{j}$, for all $j\in[T]$
$\begin{split}\lim_{m,T\rightarrow\infty}\frac{1}{T-1}\mathbb{E}_{S_{1},\ldots,S_{T}}\sum_{i=1}^{T-1}\mathcal{L}(\hat{Q}^{\lambda_{T}}_{1:T},\mathcal{D}_{i})&\leq\lim_{m,T\rightarrow\infty}\frac{1}{T-1}\sum_{i=1}^{T-1}\mathcal{L}(h^{*}_{1:T/i},\mathcal{D}_{i}).\end{split}$ (5)
The proof mainly uses Laplace’s method (Shun & McCullagh, 1995). This result implies that (asymptotically) forgetting for a given task can be low if hypotheses that work well for other tasks can be applied to it with low loss. This may be the case due to task similarity or due to the expressive power of $\mathcal{H}$. We note that the assumption of a unique global minimum may be relaxed to a finite set of global minima by replacing the loss in the r.h.s. of (5) with the highest loss among optimal hypotheses for other tasks, i.e. $\max_{h\in\mathcal{H}:h\in\arg\min_{h}\sum_{t=1,t\neq i}^{T}\mathcal{L}(h,\mathcal{D}_{t})}\mathcal{L}(h,\mathcal{D}_{i}).$
4.2 Oracle bounds for related tasks
While the bound in Theorem 4.3 is applicable for any task distribution, tighter bounds can be derived if tasks are sufficiently similar. To that end, we define the empirical task covariance
Unlike the two-task version of the loss covariance used in Corollary 4.2, this definition compares the loss on task $i$ to the total loss for all other tasks. Intuitively, this covariance is positive if hypotheses that perform well for task $i$ tend to also work well on average for all other tasks (and vice versa for hypotheses with high loss). Since the prior $P$ is data-free w.r.t. all tasks, this is a measure of both task similarity and the expressive power of hypotheses in $\mathcal{H}$.
Using Laplace’s method as well as this covariance term yields the following improved asymptotic bound:
@@ 4.2 Oracle bounds for related tasks
[…]
$\begin{split}\lim_{m\rightarrow\infty}\frac{1}{T-1}\mathbb{E}_{S_{1},\ldots,S_{T}}\sum_{i=1}^{T-1}\mathcal{L}(\hat{Q}^{\sqrt{m}}_{1:T},\mathcal{D}_{i})&\leq\frac{1}{T-1}\sum_{i=1}^{T-1}\mathcal{L}(h^{*}_{i},\mathcal{D}_{i})+\lim_{m\rightarrow\infty}\frac{\mathbb{E}_{S}\sum_{i=1}^{T-1}\log\det\hat{\mathcal{L}}_{i}^{\prime\prime}(h^{*}_{i})}{2\sqrt{m}(T-1)}.\end{split}$ (6)
The proof is deferred to Appendix B. Assuming that the determinant of the Hessian matrix is of order less than $e^{\sqrt{m}}$, which is likely unless $h^{*}_{i}$ is a very sharp global minimum, this result shows that the task covariance condition is sufficient to ensure that no forgetting occurs. We note that since the task covariance is defined w.r.t. $h\sim P$, we can expect the expressive power of $\mathcal{H}$ to impact whether or not the assumption $\mathrm{cov}_{P}(i,[T])\geq 0$ holds. This result differs from Theorem 4.3 as the r.h.s. contains the optimal hypotheses for each task $h^{*}_{1},\ldots,h^{*}_{T}$ as opposed to hypotheses that are optimal for all other tasks. Since $h^{*}_{i}=\arg\min_{h}\mathcal{L}(h,\mathcal{D}_{i})$, Theorem 4.4 is, by definition, optimally tight for $m\rightarrow\infty$ assuming that the determinant is low.
We additionally show in Appendix B that the following non-asymptotic oracle bound holds.
@@
$\mathrm{(Forgetting,Generalization)}$ For any $\lambda_{T}>0$, assuming all $Q_{1:j}$ are empirical Gibbs posteriors, and that $\mathrm{cov}_{P}(i,[T])\geq 0$, for any sample of training sets $S_{j}\sim\mathcal{D}_{j}$, then $\forall i\in[T]$
@@
Theorem 4.5 is somewhat surprising, as it implies a sufficient condition for learning without forgetting that, for each individual task, does not become worse with the number of tasks and has no direct dependence on task order or on the length of time a task has not been seen. While the r.h.s. in (7) contains a constant term $\mathcal{L}(P,\mathcal{D}_{i})$, this term does not depend on the number of total tasks $T$.
Although the resulting upper bound is potentially quite loose, it can be useful if we assume that the prior $P$ is some pre-trained model that already achieves low loss on all tasks. In this case, the Gibbs posterior allows us to iteratively fine-tune the pre-trained model while guaranteeing that the backward transfer loss is no worse than the performance of the original pre-trained model.